A U.S. firm is expecting cash flows of 90 million Mexican pesos and 100 million Indian rupees. The current spot exchange rates are $1 = 11.756 pesos and $1 = 45.221 rupees. Assume these cash flows are not received for one year and the expected spot rates at that time will be $1 = 11.152 pesos and $1 = 44.092 rupees. What is the difference in dollars received that was caused by the delay
Answer #1

Value of Cash flows in USD today:

Convert 90M Mexican Pesos = 90 M / 11.756 = USD 7.6557 M

COmvert 100 M Rupees = 100 M / 45.221 = USD 2.2114 M

Value of Cash flows today = USD 9.8671 M

Value of Cash flows after 1 Year:

Convert 90M Mexican Pesos = 90 M / 11.152 = USD 8.0703 M

COmvert 100 M Rupees = 100 M / 44.092 = USD 2.2680 M

Value of Cash flows after 1 Year = USD 10.3383 M

DIfference due to delay = Vakue after 1 year - Value today

= USD 10.3383 M - USD 9.8671 M

= USD 0.4712M

Pls comment, if any further assistance is required.